OpenStudy (anonymous):

I'm still having trouble with this can someone please walk me through the steps please Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -3x + 4y = 1 -6x + 6y = -3 I'm still having trouble with this can someone please walk me through the steps please Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -3x + 4y = 1 -6x + 6y = -3 @Mathematics

OpenStudy (agreene):

-3x+4y=1 -6x+6y=-3 pick one to start with, I'll start with the bottom. -6x+6y=-3 solve for one variable, I'll solve for y 6y=-3+6x y=(-3+6x)/6 y=(2x-1)/6 <= plug this in for y in the 1st eqn -3x+4y=1 -3x+4((2x-1)/6)=1 distribute the 4 -3x+4x-2=1 combine likes and add 2 to both sides x=3 <= plug in to the y= eqn to find y y=(2x-1)/6 y=(2(3)-1)/6 y=5/6 answers are: x=3 y=5/6

OpenStudy (anonymous):

huh